OpenStudy (anonymous):

Given vectors v(1)=(2,1,2) and v(2)=(0,-1,1), decide whether the subspace is a line or a plane through the origin. If it is a plane, compute a vector N that is perpendicular to that plane?

OpenStudy (anonymous):

@helder_edwin if you have time, could you help?

OpenStudy (helder_edwin):

u have decide wether the vectors are linearly independent or not.

OpenStudy (anonymous):

okay... I put them into a matrix and reduced it to see. Does that tell me if it is a line or a plane?

OpenStudy (helder_edwin):

yes. when u put them as rows and reduced the matrix the resulting matrix will tell u wether the original vector were linearly independent or not.

OpenStudy (helder_edwin):

if have two vectors left in the row-reduced matrix then the vector are independent.

OpenStudy (anonymous):

for some reason I thought you had to put them in as column vectors, but I still got they were independent. so if there are two vectors then it is a plane, if it is one vector they are dependent and a line?

OpenStudy (helder_edwin):

yes. it really doesn't matter if u put them as columns or rows. the number of pivots will be the same.

OpenStudy (anonymous):

so then once I found out it was independent, then how do I calculate a perpendicular vector? Can I use either vector as my X, because I am trying to solve the equation for N \[(X-X _{0}) N=0\] correct?

OpenStudy (helder_edwin):

the easy way is using the cross product from several variable calculus.

OpenStudy (helder_edwin):

do u know this?

OpenStudy (anonymous):

i do not remember the cross product

OpenStudy (helder_edwin):

the vector u r looking for would be \[ \large u=\begin{vmatrix} \vec{i} & \vec{j} & \vec{k}\\ 2 & 1 & 2\\ 0 & -1 & 1 \end{vmatrix} \]

OpenStudy (helder_edwin):

\[ \large u=(3,-2,-2) \]

OpenStudy (anonymous):

jyes that is what i got too, but i can figure out why the book would divide by -2, and have the answer (-3/2, 1, 1)

OpenStudy (helder_edwin):

u can use the one u and i got or any other parallel vector \[ \large (-1/2)u=(-3/2,1,1) \]

OpenStudy (anonymous):

so there is no particular reason why they would use that instead? or is there a different why the perpendicular vector could be compute that would get you that instead?

OpenStudy (helder_edwin):

yes. there's no particular reason to use any of them. for instance, i was taught to avoid (when possible) vectors with fractions. but, again, it is a matter of taste.

OpenStudy (anonymous):

thank you again for your expertise, i appreciate all your help.

OpenStudy (helder_edwin):

u r welcome see u soon